# 7491, Trondheim, Norway

### v1

@@ 5 Application to vector optimization without generalized convexity
Let $K$ be a nonempty, closed and convex subset of $\operatorname{\mathbb{R}}^{n}$ and $f=(f_{1},\dots,f_{m})$ be a continuous mapping from $K$ to $\operatorname{\mathbb{R}}^{m}$. The vector minimization problem with the feasible set $K$, vector objective function $f$ and nonnegative orthant +$\operatorname{\mathbb{R}}^{m}_{+}$ as ordering cone is written formally as follows:
@@
If $f$ is a scalar function, so $f:K\to\operatorname{\mathbb{R}}$, than we point out the fact, that in this case the problem $\operatorname{VOP}(K,f)$ is the well known scalar minimization problem, by writing it as $\operatorname{OP}(K,f)$. In this case the (weak) Pareto solutions are the minimizers of $f$ over $K$.
@@ 5.1 Two solution mappings
Then its relative interior $\operatorname{ri}\Delta$ is the set of all $\xi$ in $\Delta$ whose coordinates are positive. Clearly, $\Delta$ is nonempty, convex and compact. For each $\xi$ in $\Delta$, we define
@@
The mapping $\Theta_{1}$ is an effective tool to investigate the weak Pareto solution set and Pareto solution set of a vector optimization problem because of the following results which can be found in, for example, (ehrgott2005multicriteria, , Proposition 3.9) or (ansari2018vector, , Corollaries 4.1–4.3):
@@
[…]
@@
For $m=1$ this is evident, so from now on we assume $m>1$. To prove (5.5), we first prove $\operatorname{Sol^{w}}(K,f)\subseteq\Theta_{2}(\Delta)$. For this, we consider any $x\in\operatorname{Sol^{w}}(K,f)$. If there exists an index $k\in\{1,\dots,m\}$ with $f_{k}(x)=0$ then $x\in\operatorname{Sol}(K,f^{\xi})$, where $\xi=e^{k}$ is the $k$-th standard unit vector of $\operatorname{\mathbb{R}}^{m}$. Indeed, for any $y\in K$ one has $f^{\xi}(y)=\max\{0,f_{k}(y)\}\geq 0$. It is clear that $f^{\xi}(x)=0$, this implies $x\in\operatorname{Sol}(K,f^{\xi})$. From now on we assume $f(x)\in\operatorname{int}\operatorname{\mathbb{R}}^{m}_{+}$ and define $\xi\in\operatorname{\mathbb{R}}^{m}$ by
@@
Then it is $\xi\in\operatorname{ri}\Delta$ and we get $f^{\xi}(x)=\frac{1}{1/f_{1}(x)+\dots+1/f_{m}(x)}$. One has $\xi_{i}f_{i}(x)=\frac{1}{1/f_{1}(x)+\dots+1/f_{m}(x)}$, for all $i\in\{1,\dots,m\}$. Because of $x\in\operatorname{Sol^{w}}(K,f)$, for any $y\in K$, there exists an index $k\in\{1,\dots,m\}$ with $f_{k}(y)\geq f_{k}(x)$ implicating
@@
We now prove $\Theta_{2}(\Delta)\subseteq\operatorname{Sol^{w}}(K,f)$. Remark 4 points out that $\Theta_{2}(\operatorname{ri}\Delta)\subseteq\operatorname{Sol^{w}}(K,f)$. Thus, we need only to show $\Theta_{2}(\partial_{r}\Delta)\subseteq\operatorname{Sol^{w}}(K,f)$. Consider $x\in\Theta_{2}(\partial_{r}\Delta)$. Again, there exists $\xi\in\partial_{r}\Delta$ such that $x\in\operatorname{Sol}(K,f^{\xi})$. If there exists an index $k\in\{1,\dots,m\}$ with $f_{k}(x)=0$ then because of the lower boundedness we immediately get $x\in\operatorname{Sol^{w}}(K,f)$, so that from now on we can assume $f(x)\in\operatorname{int}\operatorname{\mathbb{R}}^{m}_{+}$. Define
@@
In (warburton1983quasiconcave, , p. 543) it is stated that the claim of (warburton1983quasiconcave, , Theorem 3.4) remains correct for $f$ bounded from below by any value $\gamma$. However, we want to point out that the bound zero is of utmost importance and cannot be generalized without further assumptions. In the following, we can see a counter-example.
@@
Figure 3: The relationship $\Theta_{2}(\operatorname{ri}\Delta)\varsubsetneq\operatorname{Sol^{w}}(K,f)\varsubsetneq\Theta_{2}(\operatorname{ri}\Delta)$.
@@ 5.2 Problems without imposing any generalized convexity
Assume that the domain of the solution mapping $\Theta_{1}$ given in (5.1) is non-closed. Then, the weak Pareto solution set is unbounded. Moreover, the Pareto solution set is unbounded provided that $\Theta_{1}(\operatorname{ri}\Delta)$ is nonempty.
@@
According to Remark 3, the solution mapping $\Theta_{1}$ has closed graph. The conclusions are obtained straightforwardly by applying the conclusions in Theorem 3.1 and inclusions in (5.2). ∎
@@
Consider the bicriteria optimization problem $\operatorname{VOP}(K,f)$ in $\operatorname{\mathbb{R}}^{2}$, with the feasible $K=\operatorname{\mathbb{R}}^{2}$ and two objective functions given by
@@
One can see that the functions are not bounded from below on $K$. Thus, two initial points $(0,1),(1,0)$ in $\Delta$ do not belong to the domain of the solution mapping $\Theta_{1}$ of $\operatorname{VOP}(K,f)$. Let $\xi\in\operatorname{ri}\Delta$, equivalently $\xi_{1}\in(0,1)$. One has $f_{\xi}=\xi_{1}f_{1}+(1-\xi_{1})f_{2}$ and an easy computation shows that
@@
where $h$ is a function depended only on $\xi_{1}$. Since the first two terms in $f_{\xi}$ are nonnegative, $f_{\xi}$ reaches its minimum if and only if these terms are zero. Hence, the solution mapping of $\operatorname{VOP}(K,f)$ is as given in (3.1) in Example 1 and its domain is
@@ 6 Application to quasi-convex vector optimization
for all $t\in(0,1)$. Equivalently, $g$ is quasi-convex if and only if the level set $[g\leq\lambda]$ is a convex set for all $\lambda\in\operatorname{\mathbb{R}}$. Hence, the solution set of the optimization problem $\operatorname{OP}(K,g)$ is convex. In addition, when the inequality in (6.1) is replaced by strict one, we say that $g$ is strictly quasi-convex on $K$. Recall that if $g$ is strictly quasi-convex then the solution set of $\operatorname{OP}(K,g)$ has at most one element.
@@ 6.1 Stability of solutions in parametric optimization under quasi-convexity
The following lemma about stability of solutions in quasi-convex optimization plays an important role in the proof of the main result, see Theorem 6.1; Its proof can be found in (bank1983non, , Theorem 4.3.3):
@@
Assume that, for any parameter $t\in T$, $g(\cdot,t)$ is quasi-convex on $K$. Then, the solution mapping $\Theta$ is upper semicontinuous at all points $t\in T$ such that $\Theta(t)$ is nonempty and bounded.
Considering the above parametric problem under strict quasi-convexity condition, we have the following lemma which will be used in the proof of Corollary 1.
@@
The single-valuedness of $\Theta$ is implied from the strictly quasi-convexity of $g(\cdot,t)$ on $K$. Hence, for arbitrary $t\in\operatorname{dom}\Theta$, $\Theta(t)$ is the unique isolated local solution of $\operatorname{OP}(K,g(\cdot,t))$. Applying (klatte1985stability, , Theorem 1) to the considered parametric optimization problem, we conclude that there exists a bounded and open set $V\supseteq\Theta(t)$ and a neighborhood $U$ of $t$ such that, for all $t^{\prime}\in U$, $\Theta(t^{\prime})$ is nonempty and contained in $V$. Hence, the domain of $\Theta$ is open and $\Theta$ is continuous on this set. ∎
@@ 6.2 Problems governed by quasi-convex functions
Under quasi-convexity, to ensure disconnectedness of all components of the weak Pareto solution set, the authors in (hoa2007unbounded, , Theorem 3.2) used a condition which says that the number of components of the set is finite. In the following theorem, we propose two other conditions concerning the solution mapping $\Theta_{2}$ given in (5.4) to replace it. Let
$\mathcal{O}=\{\xi\in\operatorname{dom}\Theta_{2}:\Theta_{2}(\xi)\text{ is bounded}\}.$ (6.3)
@@
$\mathcal{O}$ is open in $\Delta$.
If $\operatorname{Sol^{w}}(K,f)$ is disconnected, then all the components are unbounded. Consequently, if the set is bounded then it is connected.
@@
Since the objective functions are bounded from below by 00 on $K$, thanks to Theorem 5.1, $\operatorname{Sol^{w}}(K,f)=\Theta_{2}(\Delta)$. Moreover, $f_{1},\dots,f_{m}$ are quasi-convex, from (ansari2013generalized, , Theorem 1.31), $f^{\xi}$ given in (5.3) is also quasi-convex on $K$, for $\xi\in\Delta$, and then $\Theta_{2}$ has convex values.
According to Lemma 5, $\Theta_{2}$ is upper semi-continuous on $\mathcal{O}$. Therefore, the desired conclusion is obtained by applying directly Theorems 4.1 and 4.2 for $\Theta_{2}$ with the note that $\Theta_{2}$ has closed graph and connected values. ∎
@@
Let all $f_{i}$ be quasi-convex and strictly bounded from below by 00 on $K$. Suppose that $\Theta_{2}$ is single-valued on its domain. If $\operatorname{Sol}(K,f)$ is disconnected then all its components are unbounded.
@@
From the assumption, the domain of $\Theta_{2}$ is $\mathcal{O}$, where $\mathcal{O}$ is defined by (6.3). Hence, $\mathcal{O}$ is the set of $\xi$ in $\Delta$ such that $\Theta_{2}(\xi)$ has only one element. Thanks to (warburton1983quasiconcave, , Theorem 3.5), we claim that $\operatorname{Sol}(K,f)$ coincides with $\operatorname{Sol^{w}}(K,f)$.
@@
$\Theta_{2}(\xi_{1})=\left\{\begin{array}[]{cl}\emptyset&\ \ \hbox{ if }\ 0\leq\xi_{1}<\frac{1}{2}\hbox{ or }\ \frac{5}{6}<\xi_{1}\leq 1,\\ \left\{\Big{(}\frac{1+2\xi_{1}}{1-2\xi_{1}},0\Big{)}\right\}&\ \ \hbox{ if }\qquad\quad\ \frac{1}{6}\leq\xi_{1}<\frac{1}{2},\\ \left\{\Big{(}\frac{3-2\xi_{1}}{2\xi_{1}-1},4\Big{)}\right\}&\ \ \hbox{ if }\qquad\quad\ \frac{1}{2}<\xi_{1}\leq\frac{5}{6}.\end{array}\right.$
Clearly, the domain of $\Theta_{2}$ is the union of two open intervals $\left(\frac{1}{6},\frac{1}{2}\right)\cup\left(\frac{1}{2},\frac{5}{6}\right)$; it is disconnected. The mapping is single-valued on its domain. According to Corollary 1, the Pareto solution set and weak Pareto solution set are coincident. They are the union of two disconnected parallel lines:

### v2

@@ 5 Application to vector optimization without generalized convexity
Let $K$ be a nonempty and closed subset of $\operatorname{\mathbb{R}}^{n}$ and $f=(f_{1},\dots,f_{m})$ be a continuous mapping from $K$ to $\operatorname{\mathbb{R}}^{m}$. The vector minimization problem with the feasible set $K$, vector objective function $f$ and nonnegative orthant +$\operatorname{\mathbb{R}}^{m}_{+}$ as ordering cone is written formally as follows:
@@
If $f$ is a scalar function, so $f:K\to\operatorname{\mathbb{R}}$, then we point out the fact, that in this case the problem $\operatorname{VOP}(K,f)$ is the well known scalar minimization problem, by writing it as $\operatorname{OP}(K,f)$. In this case the (weak) Pareto solutions are the minimizers of $f$ over $K$.
@@ 5.1 Two solution mappings
Then, its relative interior $\operatorname{ri}\Delta$ is the set of all $\xi$ in $\Delta$ whose coordinates are positive. Clearly, $\Delta$ is nonempty, convex and compact. For each $\xi$ in $\Delta$, we define
@@
The mapping $\Theta_{1}$ is an effective tool to investigate the weak Pareto solution set and Pareto solution set of a vector optimization problem because of the following results which can be found in, for example, (ehrgott2005multicriteria, , Proposition 3.9) or (ansari2018vector, , Corollaries 4.1, 4.2 & 4.5):
@@
$\Theta_{1}(\Delta)\subseteq\operatorname{Sol^{w}}(K,f)\ \text{and}\ \ \Theta_{1}(\operatorname{ri}\Delta)\subseteq\operatorname{Sol}(K,f).$ (5.2)
Furthermore, if $K$ is convex and $f$ is convex, then $\Theta_{1}(\Delta)=\operatorname{Sol^{w}}(K,f)$.
@@
For $m=1$ this is evident, so from now on we assume $m>1$. To prove (5.5), we first prove $\operatorname{Sol^{w}}(K,f)\subseteq\Theta_{2}(\Delta)$. For this, we consider any $x\in\operatorname{Sol^{w}}(K,f)$. If there exists an index $k\in\{1,\dots,m\}$ with $f_{k}(x)=0$, then $x\in\operatorname{Sol}(K,f^{\xi})$, where $\xi=e^{k}$ is the $k$-th standard unit vector of $\operatorname{\mathbb{R}}^{m}$. Indeed, for any $y\in K$ one has $f^{\xi}(y)=\max\{0,f_{k}(y)\}\geq 0$. It is clear that $f^{\xi}(x)=0$, which implies $x\in\operatorname{Sol}(K,f^{\xi})$. From now on we assume $f(x)\in\operatorname{int}\operatorname{\mathbb{R}}^{m}_{+}$ and define $\xi\in\operatorname{\mathbb{R}}^{m}$ by
@@
Then, it is $\xi\in\operatorname{ri}\Delta$ and we get
$f^{\xi}(x)=\frac{1}{1/f_{1}(x)+\dots+1/f_{m}(x)}.$
One has
$\xi_{i}f_{i}(x)=\frac{1}{1/f_{1}(x)+\dots+1/f_{m}(x)},$
for all $i\in\{1,\dots,m\}$. Because of $x\in\operatorname{Sol^{w}}(K,f)$, for any $y\in K$, there exists an index $k\in\{1,\dots,m\}$ with $f_{k}(y)\geq f_{k}(x)$ implicating
@@
We now prove that $\Theta_{2}(\Delta)\subseteq\operatorname{Sol^{w}}(K,f)$. Remark 4 points out that $\Theta_{2}(\operatorname{ri}\Delta)\subseteq\operatorname{Sol^{w}}(K,f)$. Thus, we only need to show $\Theta_{2}(\partial_{r}\Delta)\subseteq\operatorname{Sol^{w}}(K,f)$,
where $\partial_{r}\Delta$ denotes the relative boundary $\partial_{r}\Delta=\overline{\Delta}\setminus\operatorname{ri}\Delta$. Consider $x\in\Theta_{2}(\partial_{r}\Delta)$. Again, there exists $\xi\in\partial_{r}\Delta$ such that $x\in\operatorname{Sol}(K,f^{\xi})$. If there exists an index $k\in\{1,\dots,m\}$ with $f_{k}(x)=0$, then because of the lower boundedness we immediately get $x\in\operatorname{Sol^{w}}(K,f)$, so that from now on we can assume $f(x)\in\operatorname{int}\operatorname{\mathbb{R}}^{m}_{+}$. Define
@@
In (warburton1983quasiconcave, , p. 543) it is stated that the claim of (warburton1983quasiconcave, , Theorem 3.4) remains correct for $f$ bounded from below by any value $\gamma$. However, we want to point out that the bound zero is of utmost importance and cannot be generalized without further assumptions. In the following, we can see a counter-example in which $\operatorname{Sol^{w}}(K,f)\neq\Theta_{2}(\Delta)$.
@@
Figure 3: The relationship $\Theta_{2}(\operatorname{ri}\Delta)\varsubsetneq\operatorname{Sol^{w}}(K,f)\varsubsetneq\Theta_{2}(\Delta)$.
@@ 5.2 Problems without imposing any generalized convexity
Assume that the domain of the solution mapping $\Theta_{1}$ given in (5.1) is non-closed. Then, the weak Pareto solution set is unbounded. Moreover, the Pareto solution set is unbounded provided that $(\operatorname{ri}\Delta)\cap\operatorname{dom}\Theta_{1}$ is dense in $\operatorname{dom}\Theta_{1}$.
@@
According to Remark 3, the solution mapping $\Theta_{1}$ has closed graph. The conclusions are obtained straightforwardly by applying the conclusions from Theorem 3.1 and the inclusions in (5.2). ∎
Remark 7
@@
Consider the bicriteria optimization problem $\operatorname{VOP}(K,f)$ in $\operatorname{\mathbb{R}}^{2}$, with the feasible set $K=\operatorname{\mathbb{R}}^{2}$ and two objective functions given by
@@
One can see that the functions are not bounded from below on $K$. Thus, the two end points $(0,1),(1,0)$ in $\Delta$ do not belong to the domain of the solution mapping $\Theta_{1}$ of $\operatorname{VOP}(K,f)$. Let $\xi\in\operatorname{ri}\Delta$, equivalently $\xi_{1}\in(0,1)$. One has $f_{\xi}=\xi_{1}f_{1}+(1-\xi_{1})f_{2}$ and an easy computation shows that
@@
where $h$ is a function depended only on $\xi_{1}$. Since the first two terms in $f_{\xi}$ are nonnegative, $f_{\xi}$ reaches its minimum if and only if these terms are zero. Hence, the solution mapping $\Theta_{1}$ of $\operatorname{VOP}(K,f)$ is as given in (3.1) in Example 1 and its domain is
@@
By repeating the argument in the proof of Theorem 5.2, we have the same conclusion for the mapping $\Theta_{2}$ given in (5.4). If the objective functions are bounded from below by 00 and the domain of the solution mapping $\Theta_{2}$ given in (5.4) is non-closed, then the weak Pareto solution set is unbounded.
@@ 6 Application to quasi-convex vector optimization
for all $t\in(0,1)$. Equivalently, $g$ is quasi-convex if and only if the level set $[g\leq\lambda]$ is a convex set for all $\lambda\in\operatorname{\mathbb{R}}$. Hence, the solution set of the optimization problem $\operatorname{OP}(K,g)$ is convex. In addition, when the inequality in (6.1) is replaced by the strict one, we say that $g$ is strictly quasi-convex on $K$. Recall that if $g$ is strictly quasi-convex, then the solution set of $\operatorname{OP}(K,g)$ has at most one element.
@@ 6.1 Stability of solutions in parametric optimization under quasi-convexity
The following lemma about stability of solutions in quasi-convex optimization plays an important role in this section; its proof can be found in (bank1983non, , Theorem 4.3.3):
@@
Assume that, for any parameter $t\in T$, $g(\cdot,t)$ is quasi-convex on $K$. Then, the solution mapping $\Theta$ defined in (6.2) is upper semi-continuous at all points $t\in T$ for which $\Theta(t)$ is nonempty and bounded.
Considering the above parametric problem under the condition of strict quasi-convexity, we have the following lemma which will be used in the proof of Corollary 1.
@@
The single-valuedness of $\Theta$ is implied by the strict quasi-convexity of $g(\cdot,t)$ on $K$. Hence, for arbitrary $t\in\operatorname{dom}\Theta$, $\Theta(t)$ is the unique isolated local solution of $\operatorname{OP}(K,g(\cdot,t))$. Applying (klatte1985stability, , Theorem 1) to the considered parametric optimization problem, we conclude that there exists a bounded and open set $V\supseteq\Theta(t)$ and a neighborhood $U$ of $t$ such that, for all $t^{\prime}\in U$, $\Theta(t^{\prime})$ is nonempty and contained in $V$. Hence, the domain of $\Theta$ is open and $\Theta$ is continuous on this set. ∎
The above lemmas will be used in the proofs of Theorem 6.1, Theorem 6.2, and Theorem 6.3 where the parameter space $T$ is $\Delta$ and the solution mapping $\Theta$ is $\Theta_{1}$ or $\Theta_{2}$.
@@ 6.2 Problems governed by quasi-convex functions
To ensure unboundedness of all components of the weak Pareto solution set of a quasi-convex vector optimization problem, the authors in (hoa2007unbounded, , Theorem 3.2) used the condition that the number of components of the solution set is finite. In the following theorem, we propose two other conditions concerning the solution mapping $\Theta_{2}$ given in (5.4) to replace it. Let
$\mathcal{O}_{2}=\{\xi\in\operatorname{dom}\Theta_{2}:\Theta_{2}(\xi)\text{ is bounded}\}.$ (6.3)
@@
$\mathcal{O}_{2}$ is open in $\Delta$.
If $\operatorname{Sol^{w}}(K,f)$ is disconnected, then all its components are unbounded. Consequently, if the set is bounded then it is connected.
@@
Since the objective functions are bounded from below by 00 on $K$, thanks to Theorem 5.1, $\operatorname{Sol^{w}}(K,f)=\Theta_{2}(\Delta)$. Moreover, $f_{1},\dots,f_{m}$ are quasi-convex, from (ansari2013generalized, , Theorem 1.31), $f^{\xi}$ given in (5.3) is also quasi-convex on $K$, for $\xi\in\Delta$, and, then $\Theta_{2}$ has convex values.
According to Lemma 5, $\Theta_{2}$ is upper semi-continuous on $\mathcal{O}_{2}$. Therefore, the desired conclusion is obtained by applying directly Theorems 4.1 and 4.2 for $\Theta_{2}$ with the note that $\Theta_{2}$ has closed graph and connected values. ∎
@@
Let all $f_{i}$ be quasi-convex and strictly bounded from below by 00 on $K$. Suppose that $\Theta_{2}$ is single-valued on its domain. If $\operatorname{Sol}(K,f)$ is disconnected, then all its components are unbounded.
@@
From the assumption, the domain of $\Theta_{2}$ is $\mathcal{O}_{2}$, where $\mathcal{O}_{2}$ is defined by (6.3). Hence, $\mathcal{O}_{2}$ is the set of $\xi$ in $\Delta$ such that $\Theta_{2}(\xi)$ has only one element. Thanks to (warburton1983quasiconcave, , Theorem 3.5), we claim that $\operatorname{Sol}(K,f)$ coincides with $\operatorname{Sol^{w}}(K,f)$.
@@
[…]
Clearly, the domain of $\Theta_{2}$ is the union of two open intervals $\left(\frac{1}{6},\frac{1}{2}\right)\cup\left(\frac{1}{2},\frac{5}{6}\right)$; it is disconnected. The mapping is single-valued on its domain. According to Corollary 1, the Pareto solution set and weak Pareto solution set coincide. They are the union of two disconnected parallel rays:
@@
In the case that all objective functions are strictly quasi-convex, $\Theta_{2}$ is single-valued on its domain. Therefore, we have the following result:
Corollary 2
Suppose that all $f_{i}$ are strictly quasi-convex and strictly bounded from below by 00 on $K$. Then, one has $\operatorname{Sol}(K,f)=\operatorname{Sol^{w}}(K,f)$. Moreover, if $\operatorname{Sol}(K,f)$ is disconnected, then all its components are unbounded.
Proof
For $\xi\in\Delta$, $f^{\xi}$ is also strictly quasi-convex on $K$ (see (warburton1983quasiconcave, , Lemma 4.2)); Hence $\Theta_{2}$ is single-valued on its domain. The conclusions are implied straightforwardly by Corollary 1. ∎
In the following theorem, we provide a necessary and sufficient condition for unboundedness of the weak Pareto solution set of $\operatorname{VOP}(K,f)$ under certain conditions:
Theorem 6.2
Let $f_{1},\dots,f_{m}$ be quasi-convex and bounded from below by 00 on $K$. Assume that $\Theta_{2}$ has bounded values. Then, the weak Pareto solution set is unbounded if and only if $\operatorname{dom}\Theta_{2}$ is non-closed in $\Delta$.
Proof
According to Theorem 5.1, one has $\operatorname{Sol^{w}}(K,f)=\Theta_{2}(\Delta)$. Clearly, if $\operatorname{dom}\Theta_{2}$ is non-closed in $\Delta$, thanks to Remark 8, the weak Pareto solution set is unbounded.
We now prove the inverse conclusion. Assume that $\operatorname{Sol^{w}}(K,f)$ is unbounded, but on the contrary, $\operatorname{dom}\Theta_{2}$ is closed in $\Delta$. Let $\{x^{k}\}$ be an unbounded sequence in $\operatorname{Sol^{w}}(K,f)$. Because of $\operatorname{Sol^{w}}(K,f)=\Theta_{2}(\Delta)$, there exist $\{\xi^{k}\}\subset\operatorname{dom}\Theta_{2}\subset\Delta$ such that $x^{k}\in\operatorname{Sol}(K,f_{\xi^{k}})$, for each $k$. By boundedness and closedness of $\operatorname{dom}\Theta_{2}$, we assume that $\xi^{k}\to\bar{\xi}\in\operatorname{dom}\Theta_{2}$, i.e., $\Theta_{2}(\bar{\xi})\neq\emptyset$. Since $\Theta_{2}$ has bounded values, $\Theta_{2}(\bar{\xi})$ is bounded.
Let $O$ be an open and bounded ball containing $\Theta_{2}(\bar{\xi})$. Thanks to Lemma 5, $\Theta_{2}$ is upper semi-continuous at $\bar{\xi}$, hence there exists an open neighbourhood $\Delta_{O}$ of $\bar{\xi}$ such that $\Theta_{2}(\xi)\subset O$ for all $\xi\in\Delta_{O}$. This implies that $x^{k}\in O$ for all $k$ large enough. This is impossible since $\{x^{k}\}$ is unbounded. ∎
6.3 Problems governed by convex functions
For problems given by convex data, we also have a necessary and sufficient condition for unboundedness of the weak Pareto solution set of $\operatorname{VOP}(K,f)$ as follows:
Theorem 6.3
Let $K$ be convex and $f_{1},\dots,f_{m}$ be convex over $K$. Assume that $\Theta_{1}$ has bounded values. Then, the weak Pareto solution set is unbounded if and only if $\operatorname{dom}\Phi_{1}$ is non-closed in $\Delta$.
Proof
We consider the scalarization $f_{\xi}=\xi_{1}f_{1}+\dots+\xi_{m}f_{m}$, for $\xi\in\Delta$. Clearly, $f_{\xi}$ is also convex over $K$. Moreover, according to Lemma 4, one has $\operatorname{Sol^{w}}(K,f)=\Phi_{1}(\Delta)$. The conclusion is proved by repeating the argument in the proof of Theorem 6.2. ∎